BIGFOOT (Characterization)
ShalikaSingh - 22:02, Wednesday 30 April 2025 (4011)
Aztec_60_40_1

Marc, Shalika

After solving issues of VI, I restarted the laser current. The power at polcam is 2mW now. 

I am starting a map, just to see if everything is fine. The step size of polarizer is 15deg and map count is 6.

To do:

1. The polarizer alignment should be tuned. The light after it is not linear (most of the time). 

2. I want to add a header in the measurement file, to save data about filtering order. 

Comments related to this report
MarcEisenmann - 11:10, Friday 02 May 2025 (4015)

There was some issue in filename saving that was solved.

I also remove second from the filename to avoid too long names.

This morning I found vi stuck at error message while saving the 3rd map.

The input laser current was about 1.7A.

I used some time to check for optimized laser current.

One issue we have is that somehow the powermeter monitoring the input laser power from the reflection of a PBS is not always sensitive to power fluctuations.

This is likely due to too low power making ambient light at the same order of magnitude of laser light.

It seems that from 2A, the input laser power has similar variation as the transmitted laser power over 20mn (Pnorm = 24.38+/-0.06 and 24.35+/-0.06)

This corresponds to about 5mW transmitted and 210uW for monitoring the input laser power.